OpenStudy (anonymous):

2x^2-x=-1 use the quadratic formula to solve the equation solution set is?

OpenStudy (campbell_st):

rewrite the equation \[2x^2 -x + 1 = 0\] you have a = 2, b = -1 and c = 1 then substitute the values into the equation \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] which gives \[x = \frac{1 \pm \sqrt{1^2 - 4\times2\times1}}{2\times2}\] just simplify it for your answer

OpenStudy (campbell_st):

Ohh, I forgot...you will have a complex number in the solution

OpenStudy (anonymous):

1/4(1-i{sqrt}7)?

OpenStudy (anonymous):

\[1/4(1-i \sqrt{7})\]

OpenStudy (campbell_st):

thats 1 solution the solutions are \[x = \frac{1 +i \sqrt{7}}{4}.... and ....x = \frac{1 - i \sqrt{7}}{4}\]